A particle execute simple harmonic motion about the point x = 0. At t= 0, it has displaced y = 0.37 cm, and zero velocity. If the frequency of motion is 0.25/s. Calculate the period, amplitude, the maximum speed and the maximum acceleration.  


1
Expert's answer
2021-02-15T00:56:11-0500

(a)

"T=\dfrac{1}{f}=\dfrac{1}{0.25\ Hz}=4.0\ s."

b)

"A=0.37\ cm=3.7\cdot10^{-3}\ m."

c) Let's first find angular frequency:


"\omega=2\pi f=2\pi\cdot0.25\ Hz=1.57\ \dfrac{rad}{s}."

Then, we can write the displacement of the particle:


"x(t)=0.0037\cdot cos(1.57t)."

Let's take the derivative of the displacement with respect to "t":


"v(t)=\dfrac{d}{dt}(x(t))=-0.0037\cdot1.57sin(1.57t)=-0.0058sin(1.57t)."

The maximum speed will be when "sin(1.57t)=1." Then, we get:


"v_{max}=-0.0058\ \dfrac{m}{s}."

d) Let's take the derivative of the speed with respect to "t":


"a(t)=\dfrac{d}{dt}(v(t))=-0.0058\cdot1.57cos(1.57t)=-0.0091cos(1.57t)."

The maximum acceleration will be when "cos(1.57t)=1." Then, we get:


"a_{max}=-0.0091\ \dfrac{m}{s^2}."